OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: Answer 4a2 - 20a + 25 25b^2 - 5b + 10 16x2 - 8x + 1 49x^2 - 8x + 16

OpenStudy (anonymous):

4a² − 20a + 25 = (2a – 5)² 16x² − 8x + 1 = (4x – 1)² If ax² + bx + c is a perfect square, then the discriminant is 0. The discriminant = b² - 4ac 49x² − 8x + 16: discriminant = (-8)² - 4(49)(16) = -3072 4a² − 20a + 25: discriminant = (-20)² - 4(4)(25) = 0 25b² − 5b + 10: discriminant = (-5)² - 4(25)(10) = -975 16x² − 8x + 1: discriminant = (-8)² - 4(16)(1) = 0

OpenStudy (anonymous):

first and third.